OpenStudy (anonymous):

Kevin and Conner tracked their exercise last week. Kevin ran 4 laps around the school track, eleven laps around his neighborhood, and 5 miles on the treadmill. Conner ran twenty laps around the school track and 6 miles on the treadmill. Write and simplify an expression for the total combined distance that Kevin and Conner ran. Show your work. Use t for the distance in laps around their school track and n for the distance in laps around Kevin’s neighborhood.

OpenStudy (anonymous):

kevin : 4*(school lap) +11*(neighborhood)+5 mile=4*t+11*n+5 (in miles) connor: 20*(school lap)+6 miles=20*t+6 K+C=4*t+11*n+6+20*t+6 24*t+11*n+11 (units in miles)